# CPE-Pro: A Structure-Sensitive Deep Learning Model for Protein Representation and Origin Evaluation

### v1

@@ Abstract
Protein structures are important for understanding their functions and interactions. Currently, many protein structure prediction methods are enriching the structure database. Discriminating the origin of structures is crucial for distinguishing between experimentally resolved and computationally predicted structures, evaluating the reliability of prediction methods, and guiding downstream biological studies. Building on works in structure prediction, We developed a structure-sensitive supervised deep learning model, Crystal vs Predicted Evaluator for Protein Structure (CPE-Pro), to represent and discriminate the origin of protein structures. CPE-Pro learns the structural information of proteins and captures inter-structural differences to achieve accurate traceability on four data classes, and is expected to be extended to more. Simultaneously, we utilized Foldseek to encode protein structures into “structure-sequence” and trained a protein Structural Sequence Language Model, SSLM. Preliminary experiments demonstrated that, compared to large-scale protein language models pre-trained on vast amounts of amino acid sequences, the “structure-sequences” enables the language model to learn more informative protein features, enhancing and optimizing structural representations. We have provided the code, model weights, and all related materials on https://github.com/GouWenrui/CPE-Pro-main.git.
@@
Deep learning, Strcuture Representation, Structure-sequence, Origin prediction1 Introduction
@@ 2.1 Protein Representation Learning
Integrating Sequence and Structure. Combining protein sequence and structural information not only considers the sequential characteristics of amino acids but also reveals their spatial interactions and arrangements. In this regard, DeeFRI[30] combines LSTM and graph convolutional networks (GCNs)[31] to jointly learn complex structure-function relationships. LM-GVP[32] modifies the input variables of GVP[33] by using sequence embeddings generated by ProteinBERT as input. ESM-GearNet[34] designs various fusion methods between PLMs and GearNet to investigate the effectiveness of different fusion strategies. ProtSSN[35] combines ESM2[11] with equivariant graph neural networks (E-GNNs)[36] to extract geometric features of proteins, aiming to accurately predict biological activity and thermal stability. ProstT5[37] introduces 3Di (Three-Dimensional Indexing) alphabet used by Foldseek[38] based on the ProtT5-XL-U50 model[39] and is trained on labeling translation tasks in two modes. Additionally, SaProt[40] creates a larger structural-aware vocabulary using two types of labels and is pre-trained on a large-scale protein dataset for masked language tasks. Compared to previous work, the SSLM in CPE-Pro uses only the “structure-sequences” built from 3Di alphabet for pre-training on Swiss-Prot, without relying on amino acid labels, and effectively learns protein representations by integrating existing structural information with GVP.
@@ 3.2 Model Architecture
Next, using the 3Di alphabet as the vocabulary of structural elements and based on the Transformer architecture, we pre-train a protein structural language model, SSLM, from scratch. This aims to effectively model the “structure-sequence” of proteins. The pre-training process employs the classic masked language modeling (MLM) objective[50], predicting masked elements based on the context of the “structure-sequence”. The probability distribution for predicting a masked element is used, where is the masked structural element and ,…,,,…, are its contexts. The loss function is defined as follows:
@@
Encoder of CPE-Pro. We integrated the “structure-sequence” representations output by the pre-trained protein structural sequence language model with the embeddings of protein graphs. The SSLM can learn the sequential relationships and proximity interactions between local structural elements from the “structure-sequence”. When combined with three-dimensional topological information, this approach aims to enrich and optimize the representation of protein structures. Specifically, we combined the representations obtained from the SSLM and the graph embeddings obtained from the GVP embedding layer as follows:represents the “structure-sequence” representation, where is the length of the “structure-sequence”, and is the feature dimension of each element in the “structure-sequence”. We align the representation dimensions of with using a linear transformation layer. Afterwards, and are fused with adaptive weights, i.e.,
@@ 5 Results and Analysis
Ablation Study on Components of CPE-Pro. To validate the contribution of each component designed in CPE-Pro, we conducted five sets of ablation experiments on both the C-A and C-M tasks. These variations included removing the GVP-GNN, omitting the pre-training process of SSLM, removing the SSLM, eliminating attention in the pooling layer, and using all three components together. As shown in Table 5, each component made a positive contribution to the C-M task. Performance significantly declined when using only a single type of encoder, particularly when using the SSLM alone as discussed earlier. The CPE-Pro model, which employs a pre-trained SSLM, outperformed the non-pre-trained version, indicating that the pre-training process effectively helped the model learn the structural features embedded in the “structure-sequences”. Additionally, the application of attention-masked pooling layers also positively influenced model performance, further enhancing overall effectiveness.
Case Study: Discrimination of the Structural Origins of BLAT ECOLX and CP2C9 HUMAN. BLAT ECOLX is a -lactamase protein found in Escherichia coli that can hydrolyze the -lactam ring of -lactam antibiotics, rendering these antibiotics inactive. It plays a significant role in the study of antibiotic resistance. CP2C9 HUMAN is a human Cytochrome P450 2C9 enzyme responsible for the metabolism of various drugs, including non-steroidal anti-inflammatory drugs and anticoagulants. It plays a significant role in drug metabolism and the regulation of endogenous substances. In three structural prediction models, both proteins achieved pLDDT scores above 0.9, indicating high accuracy in structure prediction with minimal deviation from the crystal structure. We input the crystal structures and predicted structures of both proteins into CPE-Pro for origin discrimination. Figure 6 demonstrates that the model successfully and confidently predicted the origins of these structures. This result highlights the robustness of the model in assessing structural origins, even in cases of very minor structural differences.
6 Discussion

### v2

# CPE-Pro: A Structure-Sensitive Deep Learning Method for Protein Representation and Origin Evaluation

@@ Abstract
Protein structures are important for understanding their functions and interactions. Currently, many protein structure prediction methods are enriching the structure database. Discriminating the origin of structures is crucial for distinguishing between experimentally resolved and computationally predicted structures, evaluating the reliability of prediction methods, and guiding downstream biological studies. Building on works in structure prediction, We developed a structure-sensitive supervised deep learning model, Crystal vs Predicted Evaluator for Protein Structure (CPE-Pro), to represent and discriminate the origin of protein structures. CPE-Pro learns the structural information of proteins and captures inter-structural differences to achieve accurate traceability on four data classes, and is expected to be extended to more. Simultaneously, we utilized Foldseek to encode protein structures into “structure-sequence” and trained a protein Structural Sequence Language Model, SSLM. Preliminary experiments demonstrated that, compared to large-scale protein language models pre-trained on vast amounts of amino acid sequences, the “structure-sequence” enables the language model to learn more informative protein features, enhancing and optimizing structural representations. We have provided the code, model weights, and all related materials on https://github.com/GouWenrui/CPE-Pro-main.git.
@@
Deep learning, Protein representation, Structure-sequence, Origin evaluation1 Introduction
@@ 2.1 Protein Representation Learning
Integrating Sequence and Structure. Combining protein sequence and structural information not only considers the sequential characteristics of amino acids but also reveals their spatial interactions and arrangements. In this regard, DeeFRI[30] combines LSTM and graph convolutional networks (GCNs)[31] to jointly learn complex structure-function relationships. LM-GVP[32] modifies the input variables of GVP[33] by using sequence embeddings generated by ProteinBERT as input. ESM-GearNet[34] designs various fusion methods between PLMs and GearNet to investigate the effectiveness of different fusion strategies. ProtSSN[35] combines ESM2[11] with equivariant graph neural networks (E-GNNs)[36] to extract geometric features of proteins, aiming to accurately predict biological activity and thermal stability. ProstT5[37] introduces 3Di (Three-Dimensional Indexing) alphabet used by Foldseek[38] based on the ProtT5-XL-U50 model[39] and is trained on labeling translation tasks in two modes. Additionally, SaProt[40] creates a larger structural-aware vocabulary using two types of labels and is pre-trained on a large-scale protein dataset for masked language tasks. Compared to previous work, the SSLM in CPE-Pro uses only the “structure-sequence” builts from 3Di alphabet for pre-training on Swiss-Prot, without relying on amino acid labels, and effectively learns protein representations by integrating existing structural information with GVP.
@@ 3.2 Model Architecture
Next, using the 3Di alphabet as the vocabulary of structural elements and based on the Transformer architecture, we pre-train a protein structural language model, SSLM, from scratch. This aims to effectively model “structure-sequences” of proteins. The pre-training process employs the classic masked language modeling (MLM) objective[50], predicting masked elements based on the context of the “structure-sequence”. The probability distribution for predicting a masked element is used, where is the masked structural element and ,…,,,…, are its contexts. The loss function is defined as follows:
@@
Encoder of CPE-Pro. We integrated the “structure-sequence” representation output by the pre-trained protein structural sequence language model with the embedding of protein graph. The SSLM can learn the sequential relationships and proximity interactions between local structural elements from the “structure-sequence”. When combined with three-dimensional topological information, this approach aims to enrich and optimize the representation of protein structures. Specifically, we combined the representations obtained from the SSLM and the graph embeddings obtained from the GVP embedding layer as follows:represents the “structure-sequence” representation, where is the length of the “structure-sequence”, and is the feature dimension of each element in the “structure-sequence”. We align the representation dimensions of with using a linear transformation layer. Afterwards, and are fused with adaptive weights, i.e.,
@@ 5 Results and Analysis
Ablation Study on Components of CPE-Pro. To validate the contribution of each component designed in CPE-Pro, we conducted five sets of ablation experiments on both the C-A and C-M tasks. These variations included removing the GVP-GNN, omitting the pre-training process of SSLM, removing the SSLM, eliminating attention in the pooling layer, and using all three components together. As shown in Table 5, each component made a positive contribution to the C-M task. Performance significantly declined when using only a single type of encoder, particularly when using the SSLM alone as discussed earlier. The CPE-Pro model, which employs a pre-trained SSLM, outperformed the non-pre-trained version, indicating that the pre-training process effectively helped the model learn the structural features embedded in “structure-sequences”. Additionally, the application of attention-masked pooling layers also positively influenced model performance, further enhancing overall effectiveness.
Case Study: Discrimination of the Structural Origins of BLAT ECOLX and CP2C9 HUMAN. BLAT ECOLX is a -lactamase protein found in Escherichia coli that can hydrolyze the -lactam ring of -lactam antibiotics, rendering these antibiotics inactive. It plays a significant role in the study of antibiotic resistance. CP2C9 HUMAN is a human Cytochrome P450 2C9 enzyme responsible for the metabolism of various drugs, including non-steroidal anti-inflammatory drugs and anticoagulants. It plays a significant role in drug metabolism and the regulation of endogenous substances. In three structural prediction models, both proteins achieved pLDDT scores above 0.9, indicating high accuracy in structure prediction with minimal deviation from the crystal structure. We input the crystal structures and predicted structures of both proteins into CPE-Pro for origin evaluation. Figure 6 demonstrates that the model successfully and confidently predicted the origins of these structures. This result highlights the robustness of the model in assessing structural origins, even in cases of very minor structural differences.
6 Conclusion